### v1

# Binary VPN Traffic Detection Using Wavelet Features and Machine Learning

Abstract
Encrypted traffic classification faces growing challenges as encryption renders traditional deep packet inspection ineffective. This study addresses binary VPN detection, distinguishing VPN-encrypted from non-VPN traffic using wavelet transform-based features across multiple machine learning models. We analyze the impact of wavelet decomposition levels and dataset filtering on classification performance. Our results demonstrate that Random Forest (RF) achieves superior performance with an F1-score of 99%, maintaining robust accuracy even after significant dataset filtering. Neural Networks (NN) show comparable effectiveness with an F1-score of 98% when trained on wavelet level 12, while Support Vector Machines (SVM) exhibit notable sensitivity to dataset reduction, with F1-scores dropping from 90% to 85% after filtering. Comparing wavelet decomposition at levels 5 and 12, we observe improved classification performance at level 12, particularly for variable traffic types, though the marginal gains may not justify the additional computational overhead. These findings establish RF as the most reliable model for VPN traffic classification while highlighting key performance tradeoffs in feature extraction and preprocessing.
Index Terms:
Encrypted traffic classification, VPN detection, wavelet transform, network traffic analysis
1 Introduction
The widespread adoption of network encryption poses a challenge for traffic classification, as traditional deep packet inspection and port-based methods are no longer effective [CiscoVNI2017, Statista2020]. VPNs are among the most commonly used encryption tools, with nearly one-third of internet users employing VPN services to protect their online activity [GlobalWebIndex2020]. While this enhances privacy and security, it also complicates traffic analysis for network management, intrusion detection, and policy enforcement.
Recent research has explored machine learning (ML) for encrypted traffic classification. The approaches in this field can be broadly categorized into three main groups. First, traditional feature-based methods have shown promise, with techniques like Sample Entropy Fingerprint and TCP characteristics heuristics achieving high accuracy in VPN detection [9353766, hanlon2024detecting]. Second, deep learning approaches have emerged as powerful tools, with innovations like FlowPics for visual representation of flow features [10724208] and advanced architectures combining CNN with attention mechanisms [Chai2024, Seydali2024]. The third category explores novel modeling paradigms, including multi-task learning with DistilBERT [Park2024] and text-to-text transformation using large language models [Luo2024].
[…]
In this work, we focus on binary VPN detection and re-evaluate the effectiveness of wavelet transform-based features for distinguishing VPN from non-VPN traffic. We assess the performance of multiple ML models, including Random Forest (RF), Neural Networks (NN), and Support Vector Machines (SVM), and analyze the impact of dataset filtering, a common preprocessing step that significantly affects classification performance and data distribution. Unlike prior work that primarily targets application classification, our analysis provides insights into the robustness of different ML models under specific data processing variations: wavelet decomposition depth (levels 5 and 12) and flow-length filtering.
Our results show that RF achieves the highest accuracy, reaching 99% with wavelet level 12 and maintaining strong performance (98%) even after dataset filtering. NNs also perform well, with accuracy dropping slightly from 98% to 96% after filtering. SVMs experience the most significant decline, with accuracy decreasing from 90% to 85% (wavelet level 12) and 88% to 83% (wavelet level 5) after filtering. Detailed misclassification analysis of the best-performing RF model reveals that Chat traffic exhibits the highest error rate, with frequent confusion between VPN and non-VPN traffic, even after filtering. In contrast, VoIP traffic is classified with perfect accuracy, showing no misclassification in either VPN or non-VPN scenarios. While deeper wavelet decomposition (level 12 versus level 5) improves classification accuracy, particularly for variable traffic types, the modest performance gains may not justify the additional computational overhead. These findings emphasize the importance of using robust models like RF for VPN detection, especially when dataset size is limited.
[…]
2 Wavelet Transform
Wavelet transform is a powerful mathematical tool for analyzing non-stationary signals by decomposing them into components of varying frequency and localization in time. Unlike Fourier transforms, which provide only frequency-domain representation, wavelet transforms preserve both time and frequency information, making them particularly suited for analyzing complex and transient signals such as network traffic [Ahad2016]. This section outlines the fundamentals of wavelet features, their general applications, and their specific use in the analysis of IP flows for VPN detection.
2.1 Wavelet Transform Fundamentals
2.1.1 Types of Wavelet Transforms
Several variations of wavelet transforms exist, each optimized for specific applications:
  • Continuous Wavelet Transform (CWT): Provides a detailed time-frequency representation of signals by continuously scaling and shifting the wavelet function, making it ideal for analyzing non-stationary signals with high precision.
  • Discrete Wavelet Transform (DWT): Operates on discrete scales and positions, enabling efficient computation and perfect reconstruction of signals. It is particularly suited for tasks such as signal compression and noise reduction.
  • Wavelet Packet Transform (PWT): Extends DWT by further decomposing both approximation and detail coefficients, offering a finer resolution for signals with intricate frequency components [RuizdelaHermosaGonzlezCarrato2014].
Among these, DWT is preferred for its balance between analytical capability and computational efficiency.
2.1.2 Discrete Wavelet Transform Fundamentals
DWT decomposes a signal into two sets of coefficients: approximation coefficients, which capture the low-frequency components, and detail coefficients, which represent the high-frequency components of the signal [91217].
  • Approximation coefficients ($A_{j}$) are computed as follows:
    $A_{j}[k]=\sum_{n}x[n]\cdot g(2k-n),$ (1)
    where $x[n]$ is the input signal, $g[n]$ represents the low-pass filter coefficients, and $2k$ is the downsampling factor.
  • Detail coefficients ($D_{j}$) are computed as:
    $D_{j}[k]=\sum_{n}x[n]\cdot h(2k-n),$ (2)
    where $h[n]$ represents the high-pass filter coefficients.
Both approximation and detail coefficients are used to compute the wavelet-based features. To simplify the notation in feature calculations, we introduce a general notation $C_{j}[k]$, where $C_{j}[k]$ represents either $A_{J}[k]$ (approximation coefficient at the final level) or $D_{j}[k]$ (detail coefficients at intermediate levels).
2.1.3 Wavelet Transform Metrics
To analyze the coefficients $C_{j}[k]$ obtained from DWT, we compute several key metrics that characterize different aspects of the signal.
  1. 1.
    Relative Wavelet Energy Vector ($E_{j}$) represents the energy distribution across different decomposition levels:
    $E_{j}=\frac{\sum_{k}C_{j}[k]^{2}}{\sum_{j}\sum_{k}C_{j}[k]^{2}}\times 100\%.$ (3)
    Here, $\sum_{k}C_{j}[k]^{2}$ represents the energy of the coefficients at level $j$, while $\sum_{j}\sum_{k}C_{j}[k]^{2}$ represents the total energy across all levels [Amin2015]. Since this metric reflects the relative energy distribution, it satisfies the constraint $\sum_{j}E_{j}=100\%$.
  2. 2.
    Absolute Mean of Coefficients: The absolute mean ($M_{j}$) at level $j$ quantifies the average magnitude of wavelet coefficients [753747]:
    $M_{j}=\frac{1}{N_{j}}\sum_{k}|C_{j}[k]|.$ (4)
    Here, $N_{j}$ represents the number of coefficients at level $j$, and $|C_{j}[k]|$ is the absolute value of each coefficient.
  3. 3.
    Standard Deviation of Coefficients: The standard deviation ($\sigma_{j}$) at level $j$ measures the spread of coefficients [7566534]:
    $\sigma_{j}=\sqrt{\frac{1}{N_{j}}\sum_{k}\left(C_{j}[k]-\bar{C}_{j}\right)^{2}}.$ (5)
    where the mean $\bar{C}_{j}$ is given by:
    $\bar{C}_{j}=\frac{1}{N_{j}}\sum_{k}C_{j}[k].$ (6)
  4. 4.
    Shannon Entropy: The Shannon entropy ($H_{j}$) at level $j$ quantifies the information content of the wavelet coefficients [Uyar2008]:
    $H_{j}=-\sum_{k}P_{j}[k]\log_{2}P_{j}[k].$ (7)
    The probability distribution $P_{j}[k]$ is defined as:
    $P_{j}[k]=\frac{|C_{j}[k]|}{\sum_{k}|C_{j}[k]|},$ (8)
    with the constraint:
    $\sum_{k}P_{j}[k]=1.$ (9)
2.2 Wavelet Features in IP Flows
IP flows consist of sequences of packets sharing common attributes (e.g., source/destination IP addresses and ports, and protocols) within a specific time window. Wavelet-based analysis is applied to numerical sequences derived from flow attributes such as packet size.
2.2.1 Decomposition Process
The DWT decomposes these sequences into detail coefficients ($D_{j}$) and approximation coefficients ($A_{j}$) over multiple decomposition levels ($J$). The optimal number of decomposition levels is determined as:
$J=\lfloor\log_{2}(n)\rfloor,$ (10)
where $n$ is the length of the sequence.
2.2.2 Feature Vector Construction
For each decomposition level, wavelet features are computed to create comprehensive feature vectors for classification. These include:
  • Energy Features: Raw energy values ($E_{j}$) at each decomposition level.
  • Statistical Features:
    • Absolute mean ($M_{j}$).
    • Standard deviation ($\sigma_{j}$).
    • Shannon entropy ($H_{j}$).
TABLE I: Comparison of Flow Counts with Different Active Timeouts
Complete Flows Act. Timeout 41 Act. Timeout 4.96
Category non-VPN VPN $\sum$ Tab. 3 [10044382] non-VPN VPN $\sum$ Filtered Reduction (%) Sec. IV.A [10044382]
Chat 1244 57 1301 1301 5928 5751 11679 10259 12.16 10498
Com. & Cont. 13591 8 13599 13599 15961 1624 17585 1470 91.64 1675
File Transfer 16420 10 16430 16430 17588 378 17966 805 95.52 851
Streaming 1759 5 1764 1764 3334 160 3494 1841 47.31 1827
VoIP 318 299 617 617 786 437 1223 242 80.21 243
3 Methodology
This section describes our approach to VPN traffic classification. We first present the dataset and our flow metering process, followed by feature extraction methodology for the wavelet-based features. We then detail the machine learning models employed and the evaluation metrics used to assess their performance.
3.1 Dataset Description
We use the VPN/non-VPN Network Application Traffic (VNAT) [10044382] dataset for our classification task. The dataset comprises 165 pcap files (82 VPN and 83 non-VPN traffic traces) with a total size of 36.1 GB. All network traffic was captured using tcpdump and stored in PCAP format, encompassing both VPN-encrypted and unencrypted flows across 33,711 connections and approximately 272 hours of capture time.
The dataset covers traffic from 10 applications categorized into 5 main groups: Streaming, VoIP, Chat, Command & Control, and File Transfer. These categories exhibit notable variations in their distribution across different metrics. For instance, while File Transfer represents 90% of the dataset’s volume, it accounts for only 40% of connections and 5% of the total capture time. Conversely, Command & Control traffic, despite comprising merely 2% of the dataset’s size, represents 40% of all connections and 47% of the total capture duration.
3.2 Flow Processing and Analysis
We utilized NFStream [Aouini2022], a flexible network data analysis framework, to process the dataset’s PCAP files. While VNAT provides both traffic traces in PCAP format and pre-extracted statistical summaries in HDF (h5) format, we chose NFStream for its ability to seamlessly integrate plugins, particularly for calculating the Wavelet Features discussed in Section 2.
To ensure consistency with the methodology described in [10044382], we extracted flow records under two conditions: first, without applying any flow timeout to capture complete flow records, and second, with an active timeout of 41 seconds, approximating the original study’s segmentation of flows into 40.96-second intervals. In the latter case, packets were assigned to 0.01-second time bins, and intervals containing fewer than 20 packets were discarded.
Table I presents the comparative flow counts, revealing that while complete flow counts match exactly with the original study, the application of timeout parameters leads to notable variations across traffic categories:
  • Chat: 1301 complete flows in both studies; our timeout yielded 10259 flows versus original 10498 (12.16% reduction after filtering)
  • Command & Control: 13599 complete flows; timeout processing produced 1470 flows versus 1675 (91.64% reduction after filtering)
  • File Transfer: 16430 complete flows; resulted in 805 flows versus 851 (95.52% reduction after filtering)
  • Streaming: 1764 complete flows; yielded 1841 flows versus 1827 (47.31% reduction after filtering)
  • VoIP: 617 complete flows; produced 242 flows versus 243 (80.21% reduction after filtering)
These substantial differences in flow counts, despite similar filtering criteria, demonstrate the sensitivity of flow segmentation techniques to minor timing parameter variations (41s vs. 40.96s), as NFStream only supports integer timeout values for performance optimization.
The reduction in flow counts is particularly pronounced in Command & Control (91.64%), File Transfer (95.52%), and VoIP (80.21%) traffic. This dramatic reduction highlights the effect of flow segmentation on dataset composition and raises potential concerns regarding class imbalance, as we will discuss in Section 5.
Beyond differences in flow segmentation, we identified two key methodological variations that distinguish our processing from the original VNAT dataset. First, NFStream processes only the first fragment (offset zero) of fragmented IP packets, as it requires flow headers for proper stream reconstruction. This means that while our flow segmentation remains comparable, NFStream does not reconstruct fragmented packets to obtain their true size, whereas the original study performed full reassembly. This methodological difference primarily affects large UDP-based transmissions where fragmentation occurs more frequently.
Second, we observed discrepancies in how transport layer sizes were computed. The original study inconsistently applied UDP header exclusions while including TCP headers in size calculations. NFStream, in contrast, consistently accounts for TCP and UDP payload size differences, ensuring uniformity in flow size calculations. While these differences do not directly impact the extracted features, they affect dataset comparability and highlight the sensitivity of flow statistics to implementation details.
3.3 Feature Extraction
For our classification task, we extract wavelet-based features to capture the temporal characteristics of packet size distributions in network flows. We apply DWT to transport-layer packet size sequences in both forward and backward directions. Following [10044382], we remove flows with fewer than 20 packets. Our choice of decomposition level J=5 is supported by the packet count distribution in our filtered dataset, as shown in Table II. With the 25th percentile at 56 packets, J=5 (scale of $322^{5}=$ 32 via Equation 10) ensures effective capture of packet variations across the majority of flows. We also evaluate J=12 to enable direct comparison with [10044382], though at this level (scale of $40962^{12}=$ 4096), the decomposition scale significantly exceeds typical flow lengths in their dataset. The highly skewed distribution (mean $=2,599$, median $=63$) indicates that while some flows contain up to 857,900 packets, such extensive flows are rare outliers. This distribution suggests that while J=12 might capture finer-grained variations in exceptionally long flows, it might provide diminishing returns for typical traffic patterns.
TABLE II: Packet Count Statistics Before and After Filtering
Statistic Before Filtering After Filtering
Count 51947 14617
Mean 733.50 2599.00
Standard Deviation 12444.49 23357.13
Minimum 1 20
25th Percentile 2 56
Median 2 63
75th Percentile 36 73
Maximum 857900 857900
From each decomposition, we compute four key wavelet-based metrics using both the approximation coefficients Equation 1 and detail coefficients Equation 2: relative wavelet energy Equation 3, absolute mean Equation 4, standard deviation Equation 5, and Shannon entropy Equation 7. These metrics quantify the distribution of energy across frequency components, capturing structural properties of packet sequences.
At each decomposition level $j$, these four metrics are computed separately for the corresponding detail coefficients $D_{j}$. Additionally, at the final decomposition level $J$, they are also computed for the approximation coefficients $A_{J}$, leading to $4J$ features from the $J$ sets of detail coefficients and $4$ features from the final approximation coefficients.
Since we extract these features separately for both the forward and backward packet sequences, the total number of wavelet-based features is given by $F_{wavelet}=8J+8$, where $8J$ accounts for the features extracted from all $J$ levels of detail coefficients, while the additional $8$ corresponds to the features from the final approximation coefficients. This results in 48 features for $J=5$ and 104 features for $J=12$, ensuring a balance between feature diversity and computational efficiency.
TABLE III: Performance Comparison of Machine Learning Models
Model Precision % Recall % F1-score %
RF12 99 99 99
RF12_filtered 98 98 98
RF5 99 99 99
RF5_filtered 98 98 98
NN12 99 98 98
NN12_filtered 96 96 96
NN5 93 93 93
NN5_filtered 93 93 93
SVM12 90 90 90
SVM12_filtered 85 85 85
SVM5 88 88 88
SVM5_filtered 83 83 83
3.4 Classification Models
To classify VPN and non-VPN traffic, we employ three machine learning models: RF, NN, and SVM. Each model is evaluated on both full and filtered datasets, with an 80/20 train-test split and fixed random seed to ensure reproducibility.
3.4.1 Random Forest
RF is implemented as an ensemble-based classifier, leveraging decision tree aggregation to improve generalization. It is well-suited for high-dimensional feature spaces and has been widely used in network traffic classification due to its robustness and interpretability. Default hyperparameters were used, as preliminary evaluation showed competitive performance across the dataset without requiring extensive optimization.
3.4.2 Neural Networks
NN model consists of three fully connected layers: an input layer with 64 neurons, a hidden layer with 32 neurons using ReLU activation, and an output layer with a sigmoid activation function for binary classification. The model is trained using the Adam optimizer with binary cross-entropy loss for 20 epochs.
3.4.3 Support Vector Machine
SVM is trained using a linear kernel and optimized with LinearSVC, configured with a maximum iteration limit of 10,000 to ensure convergence. We apply feature scaling prior to training to stabilize optimization.
The chosen model configurations demonstrated strong classification performance across the dataset without requiring extensive hyperparameter tuning. Given their efficiency and stability, they provide the groundwork for future research employing hyperparameter optimization and alternative models.
Figure 1: Accuracy Comparison of ML Models with and without Filtering.
3.5 Performance Metrics
We evaluate model performance using standard classification metrics: accuracy, precision, recall, and F1-score. Additionally, we analyze confusion matrices and misclassified samples to assess model behavior across different traffic categories. Misclassification patterns provide insights into category-specific classification challenges and help identify traffic types that are particularly difficult to distinguish under VPN obfuscation.
Beyond classification metrics, we investigate the effect of filtering on model performance. Since filtering reduces dataset size by discarding flows with fewer than 20 packets, it may introduce biases or reduce model generalizability.The impact of filtering on different models will be discussed in Section 4.
(a) RF12
(b) RF12_filtered
(c) RF5
(d) RF5_filtered
Figure 2: Confusion Matrices for RF with 5 and 12 Wavelet Decomposition Levels
(a) RF12
(b) RF12_filtered
(c) RF5
(d) RF5_filtered
Figure 3: Distribution of Misclassified Samples Across Different Traffic Categories for RF 12 levels
4 Results
Table III summarizes our experimental results. For visual comparison, we provide the accuracy scores separately as bar plots in Fig. 1. Models are denoted with suffixes ‘5’ or ‘12’ indicating the wavelet decomposition levels used for feature extraction, while the suffix ‘_filtered’ identifies models trained on datasets where, following the methodology in [10044382], flows with fewer than 20 packets were removed.
From Table III, we observe that RF maintains consistently high F1-scores of 99% at both decomposition levels, with only a minor decrease to 98% after filtering. NNs exhibit sensitivity to decomposition levels, with NN12 achieving an F1-score of 98%, slightly decreasing to 96% after filtering, while NN5 remains stable at 93% across both filtered and unfiltered scenarios. SVMs show the most significant impact from filtering. The F1-score for SVM12 drops from 90% to 85%, while SVM5 declines from 88% to 83% after filtering, indicating that SVM remains the most affected model by dataset reduction.
The accuracy patterns in Fig. 1 align with these observations, with RF maintaining stability at both decomposition levels, showing only minimal reductions from 99% to 98% after filtering. NN demonstrates moderate sensitivity, with NN12 accuracy decreasing from 98% to 96%, while NN5 remains unchanged at 93% post-filtering. SVM exhibits the most noticeable filtering impact, with SVM12 accuracy declining from 90% to 85% and SVM5 dropping from 88% to 83%, reinforcing that SVM is the most sensitive model to dataset reduction.
Since RF achieved consistently the best performance across both 5 and 12-level decomposition configurations, we focus our detailed analysis on its confusion matrices and misclassifications.
Fig. 2 illustrates the classification performance across different RF configurations. RF12 achieves 100% accuracy for non-VPN traffic and 97% for VPN traffic, with minimal misclassification rates of 0% and 3%, respectively. The filtered version maintains strong performance, with 99% accuracy for non-VPN and 98% for VPN traffic, showing a slight reduction in misclassification rates. RF5 exhibits similar performance, achieving 100% accuracy for non-VPN and 97% for VPN traffic. After filtering, RF5 shows a small drop in non-VPN classification accuracy to 98%, while VPN accuracy improved to 98%. These results confirm RF’s robustness, with only minimal variations in misclassification rates across different wavelet levels and filtering conditions.
Fig. 3 reveals category-specific classification challenges. Chat traffic presents the highest misclassification counts, with RF12 showing 46 total misclassifications (17 non-VPN, 29 VPN) before filtering, reducing to 35 (20 non-VPN, 15 VPN) after filtering. File Transfer and Command & Control categories show marked improvement with filtering—RF12’s misclassifications in File Transfer drop from 8 to 3 cases. RF5 shows similar patterns but with generally higher misclassification counts across all categories, particularly visible in the Chat and Streaming categories, reinforcing the advantage of higher decomposition levels for accurate classification.
However, these reduced misclassification counts after filtering should be interpreted with caution. As shown in Table I, filtering dramatically reduces the available samples—up to 95.52% in File Transfer and 91.64% in Command & Control categories. Thus, the lower misclassification counts primarily reflect the reduced dataset size rather than actual performance improvement. This observation aligns with the accuracy and F1-score metrics in Table III, which show that filtering generally decreases model performance despite the apparent reduction in misclassifications.
5 Discussion
RF outperforms NN and SVM in VPN classification across all tested conditions. RF’s ensemble structure handles missing information in shorter flows effectively, while SVM degrades due to its sensitivity to feature sparsity and distribution shifts.
The superior performance of RF and NN with J=12 over J=5 highlights the benefits of finer-grained spectral features, particularly in Chat and Streaming traffic. However, RF12 shows diminishing returns over RF5, suggesting that beyond a certain threshold, deeper wavelet decomposition adds little value. Filtering disproportionately reduces File Transfer and Command & Control traffic, exacerbating class imbalance and increasing misclassification risks, especially in underrepresented categories.
Misclassification analysis shows Chat traffic has the highest error rate, likely due to overlapping encrypted and non-encrypted patterns, while VoIP remains perfectly classified by RF, indicating distinct traffic characteristics.
Beyond accuracy, these findings highlight the importance of structured feature spaces in encrypted traffic analysis. NN models’ sensitivity to dataset variations suggests benefits from pretraining or feature augmentation. Future work could explore adaptive wavelet decomposition, focusing on the most informative spectral bands to improve efficiency and scalability.
6 Comparison with VNAT Study
[…]
10044382 [10044382] introduced VNAT, focusing on uncertainty quantification and out-of-distribution (OOD) detection. While both studies use wavelet-based features, they differ in key aspects.
The primary distinction is in objectives: VNAT targets application classification using application-specific patterns, while we focus on binary VPN detection. Both studies confirm wavelet-based features’ effectiveness in encrypted traffic classification.
Our approaches to flow processing differ. VNAT uses fixed 40.96-second intervals with a 20-packet minimum, while we compare expire flows with 41-second timeouts. Similarly, VNAT analyzes filtering through byte-level reductions, whereas our flow-count approach reveals disproportionate impacts on File Transfer and Command & Control traffic, particularly affecting SVM performance.
Feature extraction methodologies are similar, but our study systematically evaluates two decomposition depths, showing that J=12 performs better than J=5, though with diminishing returns. Unlike VNAT, which combines statistical features with wavelet representations and only uses J=12, we conducted an additional experiment beyond our main wavelet-based analysis. This experiment used only carefully selected conventional features: mean, minimum, maximum, and standard deviation of inter-arrival times and packet sizes, calculated across bidirectional, forward, and backward directions. This conventional features approach achieved performance matching our wavelet-based results at both decomposition depths (J=5 and J=12).
While this finding is noteworthy, the current landscape of publicly available VPN datasets is limited. Besides ISCXVPN2016 [DraperGil2016], which has known inconsistencies [10044382], VNAT remains the only publicly accessible dataset for VPN traffic analysis to our knowledge. This highlights a critical need for new, comprehensive VPN datasets that would enable more extensive evaluation of both conventional and wavelet-based approaches across different classification scenarios.
The studies also differ in model architecture: VNAT employs a Prototypical Network for application classification and OOD detection, while we demonstrate RF’s comparable performance without specialized architectures. However, VNAT’s OOD detection capability remains advantageous for handling unseen applications.
7 Conclusion
This study evaluated binary VPN traffic classification using wavelet-based features and machine learning models, focusing on the impact of wavelet decomposition levels and dataset filtering. Our results show that RF consistently achieves the highest classification accuracy, maintaining robustness across different preprocessing conditions, while SVM suffer the most from dataset reductions. We found that deeper wavelet decompositions improve classification performance, particularly for complex traffic types, though the marginal gains may not justify the additional computational overhead. Filtering significantly alters the dataset distribution, disproportionately affecting certain traffic categories and increasing misclassification rates. Future research could explore adaptive wavelet decomposition techniques and hybrid models that combine structured and deep learning-based feature representations for enhanced classification accuracy and generalization. Additionally, creating a new VPN and non-VPN traffic dataset could address the limitations of existing datasets, ensuring more reliable evaluation and enhancing real-world applicability.
Data Availability
This work utilizes the VNAT dataset [10044382] as the primary data source. While the original raw traffic traces from VNAT are not redistributed, all processed datasets derived from these traces—including extracted network flows, filtered datasets, and wavelet-based feature representations—are made available for research purposes. These datasets can be accessed at [github-repo].
Reproducibility
All data and artifacts necessary to reproduce the results of this study, including analysis scripts, trained models, and performance metrics, are available at [github-repo]. The provided resources have been prepared to align with the methodologies detailed in this paper, ensuring full reproducibility of the machine learning experiments. By making these materials publicly accessible, we aim to facilitate further research and validation within the community.
Acknowledgement
Supported by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences. Project no. 2024-1.2.6-EUREKA-2024-00009 has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the 2024-1.2.6-EUREKA funding scheme.
[…]

### v2

@@
 This study addresses binary VPN detection, distinguishing VPN-encrypted from non-VPN traffic using wavelet transform-based features across multiple machine learning models.
-We analyze the impact of wavelet decomposition levels and dataset filtering on classification performance.
+Unlike previous studies focused on application-level classification within encrypted traffic, we specifically evaluate the fundamental task of VPN identification regardless of application type.
+We analyze the impact of wavelet decomposition levels and dataset filtering on classification performance across significantly imbalanced data, where filtering reduces some traffic categories by up to 95%.
 Our results demonstrate that Random Forest (RF) achieves superior performance with an F1-score of 99%, maintaining robust accuracy even after significant dataset filtering.
@@
 Encrypted traffic classification, VPN detection, wavelet transform, network traffic analysis
-1 Introduction
+[…]
+I Introduction
 The widespread adoption of network encryption poses a challenge for traffic classification, as traditional deep packet inspection and port-based methods are no longer effective [CiscoVNI2017, Statista2020].
@@
 We assess the performance of multiple ML models, including Random Forest (RF), Neural Networks (NN), and Support Vector Machines (SVM), and analyze the impact of dataset filtering, a common preprocessing step that significantly affects classification performance and data distribution.
+Our analysis deliberately includes both filtered and unfiltered datasets to evaluate the robustness of models across varying conditions, even with the highly imbalanced class distributions present in real-world traffic.
 Unlike prior work that primarily targets application classification, our analysis provides insights into the robustness of different ML models under specific data processing variations: wavelet decomposition depth (levels 5 and 12) and flow-length filtering.
@@
 These findings emphasize the importance of using robust models like RF for VPN detection, especially when dataset size is limited.
-[…]
-2 Wavelet Transform
-Wavelet transform is a powerful mathematical tool for analyzing non-stationary signals by decomposing them into components of varying frequency and localization in time.
-Unlike Fourier transforms, which provide only frequency-domain representation, wavelet transforms preserve both time and frequency information, making them particularly suited for analyzing complex and transient signals such as network traffic [Ahad2016].
-This section outlines the fundamentals of wavelet features, their general applications, and their specific use in the analysis of IP flows for VPN detection.
-2.1 Wavelet Transform Fundamentals
-2.1.1 Types of Wavelet Transforms
-Several variations of wavelet transforms exist, each optimized for specific applications:
-  • Continuous Wavelet Transform (CWT): Provides a detailed time-frequency representation of signals by continuously scaling and shifting the wavelet function, making it ideal for analyzing non-stationary signals with high precision.
-  • Discrete Wavelet Transform (DWT): Operates on discrete scales and positions, enabling efficient computation and perfect reconstruction of signals.
-It is particularly suited for tasks such as signal compression and noise reduction.
-  • Wavelet Packet Transform (PWT): Extends DWT by further decomposing both approximation and detail coefficients, offering a finer resolution for signals with intricate frequency components [RuizdelaHermosaGonzlezCarrato2014].
-Among these, DWT is preferred for its balance between analytical capability and computational efficiency.
-2.1.2 Discrete Wavelet Transform Fundamentals
-DWT decomposes a signal into two sets of coefficients: approximation coefficients, which capture the low-frequency components, and detail coefficients, which represent the high-frequency components of the signal [91217].
-  • Approximation coefficients ($A_{j}$) are computed as follows:
-    $A_{j}[k]=\sum_{n}x[n]\cdot g(2k-n),$ (1)
-    where $x[n]$ is the input signal, $g[n]$ represents the low-pass filter coefficients, and $2k$ is the downsampling factor.
-  • Detail coefficients ($D_{j}$) are computed as:
-    $D_{j}[k]=\sum_{n}x[n]\cdot h(2k-n),$ (2)
-    where $h[n]$ represents the high-pass filter coefficients.
-Both approximation and detail coefficients are used to compute the wavelet-based features.
-To simplify the notation in feature calculations, we introduce a general notation $C_{j}[k]$, where $C_{j}[k]$ represents either $A_{J}[k]$ (approximation coefficient at the final level) or $D_{j}[k]$ (detail coefficients at intermediate levels).
-2.1.3 Wavelet Transform Metrics
-To analyze the coefficients $C_{j}[k]$ obtained from DWT, we compute several key metrics that characterize different aspects of the signal.
-  1. 1.
-    Relative Wavelet Energy Vector ($E_{j}$) represents the energy distribution across different decomposition levels:
-    $E_{j}=\frac{\sum_{k}C_{j}[k]^{2}}{\sum_{j}\sum_{k}C_{j}[k]^{2}}\times 100\%.$ (3)
-    Here, $\sum_{k}C_{j}[k]^{2}$ represents the energy of the coefficients at level $j$, while $\sum_{j}\sum_{k}C_{j}[k]^{2}$ represents the total energy across all levels [Amin2015].
-Since this metric reflects the relative energy distribution, it satisfies the constraint $\sum_{j}E_{j}=100\%$.
-  2. 2.
-    Absolute Mean of Coefficients: The absolute mean ($M_{j}$) at level $j$ quantifies the average magnitude of wavelet coefficients [753747]:
-    $M_{j}=\frac{1}{N_{j}}\sum_{k}|C_{j}[k]|.$ (4)
-    Here, $N_{j}$ represents the number of coefficients at level $j$, and $|C_{j}[k]|$ is the absolute value of each coefficient.
-  3. 3.
-    Standard Deviation of Coefficients: The standard deviation ($\sigma_{j}$) at level $j$ measures the spread of coefficients [7566534]:
-    $\sigma_{j}=\sqrt{\frac{1}{N_{j}}\sum_{k}\left(C_{j}[k]-\bar{C}_{j}\right)^{2}}.$ (5)
-    where the mean $\bar{C}_{j}$ is given by:
-    $\bar{C}_{j}=\frac{1}{N_{j}}\sum_{k}C_{j}[k].$ (6)
-  4. 4.
-    Shannon Entropy: The Shannon entropy ($H_{j}$) at level $j$ quantifies the information content of the wavelet coefficients [Uyar2008]:
-    $H_{j}=-\sum_{k}P_{j}[k]\log_{2}P_{j}[k].$ (7)
-    The probability distribution $P_{j}[k]$ is defined as:
-    $P_{j}[k]=\frac{|C_{j}[k]|}{\sum_{k}|C_{j}[k]|},$ (8)
-    with the constraint:
-    $\sum_{k}P_{j}[k]=1.$ (9)
-2.2 Wavelet Features in IP Flows
-IP flows consist of sequences of packets sharing common attributes (e.g., source/destination IP addresses and ports, and protocols) within a specific time window.
-Wavelet-based analysis is applied to numerical sequences derived from flow attributes such as packet size.
-2.2.1 Decomposition Process
-The DWT decomposes these sequences into detail coefficients ($D_{j}$) and approximation coefficients ($A_{j}$) over multiple decomposition levels ($J$).
-The optimal number of decomposition levels is determined as:
-$J=\lfloor\log_{2}(n)\rfloor,$ (10)
-where $n$ is the length of the sequence.
-2.2.2 Feature Vector Construction
-For each decomposition level, wavelet features are computed to create comprehensive feature vectors for classification. These include:
-  • Energy Features: Raw energy values ($E_{j}$) at each decomposition level.
-  • Statistical Features:
-    • Absolute mean ($M_{j}$).
-    • Standard deviation ($\sigma_{j}$).
-    • Shannon entropy ($H_{j}$).
+The remainder of this paper is structured as follows: Section II describes our methodology, including dataset preprocessing and model selection.
+Section III presents and analyzes our results.
+Finally, Section IV concludes this paper.
 TABLE I: Comparison of Flow Counts with Different Active Timeouts
-Complete Flows Act. Timeout 41 Act. Timeout 4.96
+Complete Flows Act. Timeout 41 Act. Timeout 40.96
 Category non-VPN VPN $\sum$ Tab. 3 [10044382] non-VPN VPN $\sum$ Filtered Reduction (%) Sec. IV.A [10044382]
@@
 VoIP 318 299 617 617 786 437 1223 242 80.21 243
-3 Methodology
-This section describes our approach to VPN traffic classification.
-We first present the dataset and our flow metering process, followed by feature extraction methodology for the wavelet-based features.
-We then detail the machine learning models employed and the evaluation metrics used to assess their performance.
-3.1 Dataset Description
+II Methodology
+II-A Dataset Description
 We use the VPN/non-VPN Network Application Traffic (VNAT) [10044382] dataset for our classification task.
@@
 Conversely, Command & Control traffic, despite comprising merely 2% of the dataset’s size, represents 40% of all connections and 47% of the total capture duration.
-3.2 Flow Processing and Analysis
+II-B Flow Processing and Analysis
 We utilized NFStream [Aouini2022], a flexible network data analysis framework, to process the dataset’s PCAP files.
-While VNAT provides both traffic traces in PCAP format and pre-extracted statistical summaries in HDF (h5) format, we chose NFStream for its ability to seamlessly integrate plugins, particularly for calculating the Wavelet Features discussed in Section 2.
+While VNAT provides both traffic traces in PCAP format and pre-extracted statistical summaries in HDF (h5) format, we chose NFStream for its ability to seamlessly integrate plugins, particularly for calculating the Wavelet Features discussed in Section II-C.
 To ensure consistency with the methodology described in [10044382], we extracted flow records under two conditions: first, without applying any flow timeout to capture complete flow records, and second, with an active timeout of 41 seconds, approximating the original study’s segmentation of flows into 40.96-second intervals.
 In the latter case, packets were assigned to 0.01-second time bins, and intervals containing fewer than 20 packets were discarded.
+This dual approach enables us to evaluate the impact of flow segmentation and filtering on classification performance.
 Table I presents the comparative flow counts, revealing that while complete flow counts match exactly with the original study, the application of timeout parameters leads to notable variations across traffic categories:
@@
 The reduction in flow counts is particularly pronounced in Command & Control (91.64%), File Transfer (95.52%), and VoIP (80.21%) traffic.
-This dramatic reduction highlights the effect of flow segmentation on dataset composition and raises potential concerns regarding class imbalance, as we will discuss in Section 5.
+This dramatic reduction highlights the effect of flow segmentation on dataset composition and raises potential concerns regarding class imbalance.
 Beyond differences in flow segmentation, we identified two key methodological variations that distinguish our processing from the original VNAT dataset.
@@
 While these differences do not directly impact the extracted features, they affect dataset comparability and highlight the sensitivity of flow statistics to implementation details.
-3.3 Feature Extraction
+This filtering has a non-uniform effect on VPN and non-VPN traffic distributions, altering the class balance within categories and creating a more challenging classification scenario that better reflects real-world deployment conditions.
+II-C Wavelet Features
+Wavelet transform is a time-frequency analysis method that decomposes signals into components at different scales, making it suitable for analyzing non-stationary traffic patterns.
+Unlike Fourier transforms, wavelets preserve both time and frequency information, capturing traffic bursts and periodic behaviors effectively [Ahad2016].
+The Discrete Wavelet Transform (DWT) decomposes a signal into approximation coefficients (low-frequency components) and detail coefficients (high-frequency components) at multiple levels.
+For a network flow, the signal is the sequence of packet sizes over time [91217].
+The approximation coefficients are computed as:
+$A_{j}[k]=\sum_{n}x[n]\cdot g(2k-n)$ (1)
+where $x[n]$ is the input signal, $g[n]$ represents the low-pass filter coefficients, and $2k$ is the downsampling factor. The detail coefficients are computed as:
+$D_{j}[k]=\sum_{n}x[n]\cdot h(2k-n)$ (2)
+where $h[n]$ represents the high-pass filter coefficients.
+From these coefficients, we extract four key metrics that characterize different aspects of network traffic:
+  1. 1.
+    Relative Wavelet Energy Vector ($E_{j}$) represents the energy distribution across different decomposition levels:
+    $E_{j}=\frac{\sum_{k}C_{j}[k]^{2}}{\sum_{j}\sum_{k}C_{j}[k]^{2}}\times 100\%$ (3)
+  2. 2.
+    Absolute Mean of Coefficients ($M_{j}$) quantifies the average magnitude at level $j$  [753747]:
+    $M_{j}=\frac{1}{N_{j}}\sum_{k}|C_{j}[k]|$ (4)
+  3. 3.
+    Standard Deviation of Coefficients ($\sigma_{j}$) measures coefficient spread [7566534]:
+    $\sigma_{j}=\sqrt{\frac{1}{N_{j}}\sum_{k}\left(C_{j}[k]-\bar{C}_{j}\right)^{2}}$ (5)
+  4. 4.
+    Shannon Entropy ($H_{j}$) quantifies the information content [Shannon1948]:
+    $H_{j}=-\sum_{k}P_{j}[k]\log_{2}P_{j}[k]$ (6)
+    where $P_{j}[k]=\frac{|C_{j}[k]|}{\sum_{k}|C_{j}[k]|}$.
+The optimal number of decomposition levels is determined as $J=\lfloor\log_{2}(n)\rfloor$, where $n$ is the sequence length.
+This provides a mathematical basis for our choice of decomposition levels in feature extraction.
+II-D Feature Extraction
 For our classification task, we extract wavelet-based features to capture the temporal characteristics of packet size distributions in network flows.
@@
 Our choice of decomposition level J=5 is supported by the packet count distribution in our filtered dataset, as shown in Table II.
-With the 25th percentile at 56 packets, J=5 (scale of $322^{5}=$ 32 via Equation 10) ensures effective capture of packet variations across the majority of flows.
+With the 25th percentile at 56 packets, J=5 (scale of $322^{5}=$ 32 according to the optimal decomposition level formula $J=\lfloor\log_{2}(n)\rfloor$) ensures effective capture of packet variations across the majority of flows.
 We also evaluate J=12 to enable direct comparison with [10044382], though at this level (scale of $40962^{12}=$ 4096), the decomposition scale significantly exceeds typical flow lengths in their dataset.
@@
 Maximum 857900 857900
-From each decomposition, we compute four key wavelet-based metrics using both the approximation coefficients Equation 1 and detail coefficients Equation 2: relative wavelet energy Equation 3, absolute mean Equation 4, standard deviation Equation 5, and Shannon entropy Equation 7.
+From each decomposition, we compute four key wavelet-based metrics using both the approximation coefficients and detail coefficients: relative wavelet energy, absolute mean, standard deviation, and Shannon entropy.
 These metrics quantify the distribution of energy across frequency components, capturing structural properties of packet sequences.
@@
 This results in 48 features for $J=5$ and 104 features for $J=12$, ensuring a balance between feature diversity and computational efficiency.
-TABLE III: Performance Comparison of Machine Learning Models
+II-E Classification Models
+To classify VPN and non-VPN traffic, we employ three machine learning models: RF, NN, and SVM.
+Each model is evaluated on both full and filtered datasets, with an 80/20 train-test split and fixed random seed to ensure reproducibility.
+II-E1 Random Forest
+RF is implemented as an ensemble-based classifier, leveraging decision tree aggregation to improve generalization.
+Default hyperparameters were used, as preliminary evaluation showed competitive performance across the dataset without requiring extensive optimization.
+II-E2 Neural Networks
+NN model consists of three fully connected layers: an input layer with 64 neurons, a hidden layer with 32 neurons using ReLU activation, and an output layer with a sigmoid activation function for binary classification.
+The model is trained using the Adam optimizer with binary cross-entropy loss for 20 epochs.
+II-E3 Support Vector Machine
+SVM is trained using a linear kernel and optimized with LinearSVC, configured with a maximum iteration limit of 10,000 to ensure convergence.
+We apply feature scaling prior to training to stabilize optimization.
+The chosen model configurations demonstrated strong classification performance across the dataset without requiring extensive hyperparameter tuning.
+Given their efficiency and stability, they provide the groundwork for future research employing hyperparameter optimization and alternative models.
+II-F Performance Metrics
+We evaluate model performance using standard classification metrics: accuracy, precision, recall, and F1-score.
+Additionally, we analyze confusion matrices and misclassified samples to assess model behavior across different traffic categories.
+Misclassification patterns provide insights into category-specific classification challenges and help identify traffic types that are particularly difficult to distinguish under VPN obfuscation.
+Beyond classification metrics, we investigate the effect of filtering on model performance.
+Since filtering reduces dataset size by discarding flows with fewer than 20 packets, it may introduce biases or reduce model generalizability.
+We deliberately evaluate performance on both filtered and unfiltered datasets to determine how robust each model is to variations in flow length, an important consideration for real-world deployment where flows of all lengths will be encountered.
+Section III discusses the impact of filtering on different models.
+Analyzing both filtered and unfiltered datasets, despite the predominance of very short flows (1-2 packets) in the latter, serves two important purposes: first, it establishes a performance baseline for comparison across preprocessing techniques; second, it evaluates the models’ capability to extract meaningful signal from minimal data, which is essential in scenarios where early traffic detection is required before sufficient packets have been observed.
+III Results
+Table III summarizes our experimental results across all model configurations.
+For visual comparison, we provide the accuracy scores as bar plots in Fig.
+1.
+Models are denoted with suffixes ‘5’ or ‘12’ indicating the wavelet decomposition levels used for feature extraction, while the suffix ‘_filtered’ identifies models trained on datasets where flows with fewer than 20 packets were removed.
+TABLE III: Performance Comparison of ML Models
 Model Precision % Recall % F1-score %
@@
 SVM5_filtered 83 83 83
-3.4 Classification Models
-To classify VPN and non-VPN traffic, we employ three machine learning models: RF, NN, and SVM.
-Each model is evaluated on both full and filtered datasets, with an 80/20 train-test split and fixed random seed to ensure reproducibility.
-3.4.1 Random Forest
-RF is implemented as an ensemble-based classifier, leveraging decision tree aggregation to improve generalization.
-It is well-suited for high-dimensional feature spaces and has been widely used in network traffic classification due to its robustness and interpretability.
-Default hyperparameters were used, as preliminary evaluation showed competitive performance across the dataset without requiring extensive optimization.
-3.4.2 Neural Networks
-NN model consists of three fully connected layers: an input layer with 64 neurons, a hidden layer with 32 neurons using ReLU activation, and an output layer with a sigmoid activation function for binary classification.
-The model is trained using the Adam optimizer with binary cross-entropy loss for 20 epochs.
-3.4.3 Support Vector Machine
-SVM is trained using a linear kernel and optimized with LinearSVC, configured with a maximum iteration limit of 10,000 to ensure convergence.
-We apply feature scaling prior to training to stabilize optimization.
-The chosen model configurations demonstrated strong classification performance across the dataset without requiring extensive hyperparameter tuning.
-Given their efficiency and stability, they provide the groundwork for future research employing hyperparameter optimization and alternative models.
-Figure 1: Accuracy Comparison of ML Models with and without Filtering.
-3.5 Performance Metrics
-We evaluate model performance using standard classification metrics: accuracy, precision, recall, and F1-score.
-Additionally, we analyze confusion matrices and misclassified samples to assess model behavior across different traffic categories.
-Misclassification patterns provide insights into category-specific classification challenges and help identify traffic types that are particularly difficult to distinguish under VPN obfuscation.
-Beyond classification metrics, we investigate the effect of filtering on model performance.
-Since filtering reduces dataset size by discarding flows with fewer than 20 packets, it may introduce biases or reduce model generalizability.The impact of filtering on different models will be discussed in Section 4.
+Figure 1: Accuracy Comparison with and without Filtering
 (a) RF12
@@
 (d) RF5_filtered
-Figure 3: Distribution of Misclassified Samples Across Different Traffic Categories for RF 12 levels
-4 Results
-Table III summarizes our experimental results.
-For visual comparison, we provide the accuracy scores separately as bar plots in Fig.
-1.
-Models are denoted with suffixes ‘5’ or ‘12’ indicating the wavelet decomposition levels used for feature extraction, while the suffix ‘_filtered’ identifies models trained on datasets where, following the methodology in [10044382], flows with fewer than 20 packets were removed.
+Figure 3: Distribution of Misclassified Samples Across Different Traffic Categories for RF 5 and 12 levels
 From Table III, we observe that RF maintains consistently high F1-scores of 99% at both decomposition levels, with only a minor decrease to 98% after filtering.
+This remarkable stability across both preprocessing conditions demonstrates RF’s robust feature extraction capability even with minimal data.
 NNs exhibit sensitivity to decomposition levels, with NN12 achieving an F1-score of 98%, slightly decreasing to 96% after filtering, while NN5 remains stable at 93% across both filtered and unfiltered scenarios.
-SVMs show the most significant impact from filtering.
-The F1-score for SVM12 drops from 90% to 85%, while SVM5 declines from 88% to 83% after filtering, indicating that SVM remains the most affected model by dataset reduction.
+This 5% performance gap between NN12 and NN5 suggests that NNs benefit from the additional spectral information in higher decomposition levels.
+SVMs show the most significant impact from filtering, with F1-scores for SVM12 dropping from 90% to 85% (5% decrease), while SVM5 declines from 88% to 83% (also 5% decrease), indicating consistent sensitivity to dataset reduction regardless of decomposition level.
+This sensitivity likely stems from SVM’s reliance on support vectors at the class boundary, which become more sparse after filtering.
 The accuracy patterns in Fig.
-1 align with these observations, with RF maintaining stability at both decomposition levels, showing only minimal reductions from 99% to 98% after filtering.
+1 confirm these observations.
+RF maintains high stability at both decomposition levels, showing only minimal reductions from 99% to 98% after filtering.
 NN demonstrates moderate sensitivity, with NN12 accuracy decreasing from 98% to 96%, while NN5 remains unchanged at 93% post-filtering.
-SVM exhibits the most noticeable filtering impact, with SVM12 accuracy declining from 90% to 85% and SVM5 dropping from 88% to 83%, reinforcing that SVM is the most sensitive model to dataset reduction.
-Since RF achieved consistently the best performance across both 5 and 12-level decomposition configurations, we focus our detailed analysis on its confusion matrices and misclassifications.
+SVM exhibits the most pronounced filtering impact, with SVM12 accuracy declining from 90% to 85% and SVM5 dropping from 88% to 83%.
+Since RF achieved consistently the best performance across both decomposition configurations, we focus our detailed analysis on its confusion matrices and misclassification patterns.
 Fig.
@@
 RF12 achieves 100% accuracy for non-VPN traffic and 97% for VPN traffic, with minimal misclassification rates of 0% and 3%, respectively.
-The filtered version maintains strong performance, with 99% accuracy for non-VPN and 98% for VPN traffic, showing a slight reduction in misclassification rates.
+The filtered version maintains strong performance, with 99% accuracy for non-VPN and 98% for VPN traffic.
 RF5 exhibits similar performance, achieving 100% accuracy for non-VPN and 97% for VPN traffic.
@@
 3 reveals category-specific classification challenges.
-Chat traffic presents the highest misclassification counts, with RF12 showing 46 total misclassifications (17 non-VPN, 29 VPN) before filtering, reducing to 35 (20 non-VPN, 15 VPN) after filtering.
-File Transfer and Command & Control categories show marked improvement with filtering—RF12’s misclassifications in File Transfer drop from 8 to 3 cases.
-RF5 shows similar patterns but with generally higher misclassification counts across all categories, particularly visible in the Chat and Streaming categories, reinforcing the advantage of higher decomposition levels for accurate classification.
-However, these reduced misclassification counts after filtering should be interpreted with caution.
-As shown in Table I, filtering dramatically reduces the available samples—up to 95.52% in File Transfer and 91.64% in Command & Control categories.
-Thus, the lower misclassification counts primarily reflect the reduced dataset size rather than actual performance improvement.
-This observation aligns with the accuracy and F1-score metrics in Table III, which show that filtering generally decreases model performance despite the apparent reduction in misclassifications.
-5 Discussion
-RF outperforms NN and SVM in VPN classification across all tested conditions.
-RF’s ensemble structure handles missing information in shorter flows effectively, while SVM degrades due to its sensitivity to feature sparsity and distribution shifts.
-The superior performance of RF and NN with J=12 over J=5 highlights the benefits of finer-grained spectral features, particularly in Chat and Streaming traffic.
-However, RF12 shows diminishing returns over RF5, suggesting that beyond a certain threshold, deeper wavelet decomposition adds little value.
-Filtering disproportionately reduces File Transfer and Command & Control traffic, exacerbating class imbalance and increasing misclassification risks, especially in underrepresented categories.
-Misclassification analysis shows Chat traffic has the highest error rate, likely due to overlapping encrypted and non-encrypted patterns, while VoIP remains perfectly classified by RF, indicating distinct traffic characteristics.
-Beyond accuracy, these findings highlight the importance of structured feature spaces in encrypted traffic analysis.
-NN models’ sensitivity to dataset variations suggests benefits from pretraining or feature augmentation.
-Future work could explore adaptive wavelet decomposition, focusing on the most informative spectral bands to improve efficiency and scalability.
-6 Comparison with VNAT Study
-[…]
-10044382 [10044382] introduced VNAT, focusing on uncertainty quantification and out-of-distribution (OOD) detection. While both studies use wavelet-based features, they differ in key aspects.
-The primary distinction is in objectives: VNAT targets application classification using application-specific patterns, while we focus on binary VPN detection.
-Both studies confirm wavelet-based features’ effectiveness in encrypted traffic classification.
-Our approaches to flow processing differ.
-VNAT uses fixed 40.96-second intervals with a 20-packet minimum, while we compare expire flows with 41-second timeouts.
-Similarly, VNAT analyzes filtering through byte-level reductions, whereas our flow-count approach reveals disproportionate impacts on File Transfer and Command & Control traffic, particularly affecting SVM performance.
-Feature extraction methodologies are similar, but our study systematically evaluates two decomposition depths, showing that J=12 performs better than J=5, though with diminishing returns.
-Unlike VNAT, which combines statistical features with wavelet representations and only uses J=12, we conducted an additional experiment beyond our main wavelet-based analysis.
-This experiment used only carefully selected conventional features: mean, minimum, maximum, and standard deviation of inter-arrival times and packet sizes, calculated across bidirectional, forward, and backward directions.
-This conventional features approach achieved performance matching our wavelet-based results at both decomposition depths (J=5 and J=12).
-While this finding is noteworthy, the current landscape of publicly available VPN datasets is limited.
-Besides ISCXVPN2016 [DraperGil2016], which has known inconsistencies [10044382], VNAT remains the only publicly accessible dataset for VPN traffic analysis to our knowledge.
-This highlights a critical need for new, comprehensive VPN datasets that would enable more extensive evaluation of both conventional and wavelet-based approaches across different classification scenarios.
-The studies also differ in model architecture: VNAT employs a Prototypical Network for application classification and OOD detection, while we demonstrate RF’s comparable performance without specialized architectures.
-However, VNAT’s OOD detection capability remains advantageous for handling unseen applications.
-7 Conclusion
+Chat traffic presents the highest error counts, with RF12 showing 46 total misclassifications (17 non-VPN, 29 VPN) before filtering, reducing to 35 (20 non-VPN, 15 VPN) after filtering.
+When viewed relative to category size (1301 Chat samples), this represents approximately 3.5% of Chat samples misclassified.
+The higher error rate likely stems from Chat’s variable traffic patterns and similarities between encrypted and non-encrypted protocols.
+File Transfer and Command & Control categories show marked improvement with filtering—RF12’s misclassifications in File Transfer drop from 8 to 3 cases, suggesting that longer flows contain more distinctive VPN signatures.
+RF5 shows similar patterns but with generally higher misclassification counts across all categories, particularly visible in the Chat and Streaming categories.
+Notably, as shown in Fig.
+3, VoIP traffic achieves zero misclassifications in the RF12 model under both filtered and unfiltered conditions, indicating highly distinctive traffic patterns between VPN and non-VPN variants despite the 80% dataset reduction from filtering.
+This exceptional performance suggests that VoIP applications maintain unique temporal signatures despite encryption, likely due to their real-time streaming requirements.
+When interpreting the reduced misclassification counts after filtering, it’s important to consider the dramatic reduction in available samples—up to 95.52% in File Transfer and 91.64% in Command & Control categories, as shown in Table I.
+Thus, while absolute misclassification counts decrease after filtering, the error rates as a percentage of category size actually increase slightly in several categories (File Transfer error rate increases from 0.05% to 0.37% after filtering).
+This observation aligns with the overall F1-score patterns in Table III, confirming that filtering generally causes a slight performance degradation.
+The performance comparison between J=12 and J=5 reveals a consistent advantage for deeper wavelet decomposition, with improvements ranging from 1% (RF12 vs.
+RF5) to 5% (NN12 vs.
+NN5) in F1-scores.
+However, these relatively modest gains suggest diminishing returns that may not justify the increased computational overhead in all deployment scenarios, particularly in resource-constrained environments where model efficiency is prioritized alongside accuracy.
+IV Conclusion
 This study evaluated binary VPN traffic classification using wavelet-based features and machine learning models, focusing on the impact of wavelet decomposition levels and dataset filtering.
-Our results show that RF consistently achieves the highest classification accuracy, maintaining robustness across different preprocessing conditions, while SVM suffer the most from dataset reductions.
-We found that deeper wavelet decompositions improve classification performance, particularly for complex traffic types, though the marginal gains may not justify the additional computational overhead.
-Filtering significantly alters the dataset distribution, disproportionately affecting certain traffic categories and increasing misclassification rates.
-Future research could explore adaptive wavelet decomposition techniques and hybrid models that combine structured and deep learning-based feature representations for enhanced classification accuracy and generalization.
-Additionally, creating a new VPN and non-VPN traffic dataset could address the limitations of existing datasets, ensuring more reliable evaluation and enhancing real-world applicability.
-Data Availability
+Our results show that RF consistently achieves the highest classification accuracy, maintaining robustness across different preprocessing conditions, while SVM exhibits the greatest sensitivity to dataset reductions.
+We found that deeper wavelet decompositions improve classification performance, particularly for variable traffic types, though the marginal gains may not justify the additional computational overhead.
+Future research could explore adaptive wavelet decomposition techniques that dynamically adjust to flow characteristics and investigate ensemble approaches combining multiple decomposition levels for enhanced robustness across diverse traffic conditions.
 This work utilizes the VNAT dataset [10044382] as the primary data source.
-While the original raw traffic traces from VNAT are not redistributed, all processed datasets derived from these traces—including extracted network flows, filtered datasets, and wavelet-based feature representations—are made available for research purposes.
-These datasets can be accessed at [github-repo].
-Reproducibility
-All data and artifacts necessary to reproduce the results of this study, including analysis scripts, trained models, and performance metrics, are available at [github-repo].
-The provided resources have been prepared to align with the methodologies detailed in this paper, ensuring full reproducibility of the machine learning experiments.
-By making these materials publicly accessible, we aim to facilitate further research and validation within the community.
+All processed datasets derived from these traces—including extracted network flows, filtered datasets, and wavelet-based feature representations—along with analysis scripts, trained models, and performance metrics necessary to reproduce our results are available at [github-repo], ensuring full reproducibility of the machine learning experiments.
 Acknowledgement
 Supported by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.
-Project no.
-2024-1.2.6-EUREKA-2024-00009 has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the 2024-1.2.6-EUREKA funding scheme.
+This work was also supported by project TKP2021-NVA-02 (financed under the TKP2021-NVA scheme), implemented with support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund.
 […]

### v3

@@
 I Introduction
-The widespread adoption of network encryption poses a challenge for traffic classification, as traditional deep packet inspection and port-based methods are no longer effective [CiscoVNI2017, Statista2020].
-VPNs are among the most commonly used encryption tools, with nearly one-third of internet users employing VPN services to protect their online activity [GlobalWebIndex2020].
+The widespread adoption of network encryption poses a challenge for traffic classification, as traditional deep packet inspection and port-based methods are no longer effective [1, 2].
+VPNs are among the most commonly used encryption tools, with nearly one-third of internet users employing VPN services to protect their online activity [3].
 While this enhances privacy and security, it also complicates traffic analysis for network management, intrusion detection, and policy enforcement.
@@
 The approaches in this field can be broadly categorized into three main groups.
-First, traditional feature-based methods have shown promise, with techniques like Sample Entropy Fingerprint and TCP characteristics heuristics achieving high accuracy in VPN detection [9353766, hanlon2024detecting].
-Second, deep learning approaches have emerged as powerful tools, with innovations like FlowPics for visual representation of flow features [10724208] and advanced architectures combining CNN with attention mechanisms [Chai2024, Seydali2024].
-The third category explores novel modeling paradigms, including multi-task learning with DistilBERT [Park2024] and text-to-text transformation using large language models [Luo2024].
-[…]
+First, traditional feature-based methods have shown promise, with techniques like Sample Entropy Fingerprint and TCP characteristics heuristics achieving high accuracy in VPN detection [4, 5].
+Second, deep learning approaches have emerged as powerful tools, with innovations like FlowPics for visual representation of flow features [6] and advanced architectures combining CNN with attention mechanisms [7, 8].
+The third category explores novel modeling paradigms, including multi-task learning with DistilBERT [9] and text-to-text transformation using large language models [10].
+Notably, [11][11] introduced an ML framework that combines wavelet-based features with uncertainty quantification for application-level classification across VPN and non-VPN traffic.
+However, the related but distinct problem of binary VPN detection—determining whether traffic is VPN-encrypted regardless of application—has received less attention and presents unique value for network policy enforcement and security monitoring.
+While wavelet transforms have proven effective in network anomaly detection[12, 13], their potential for binary VPN classification remains unexplored.
+Building on VNAT’s approach, our study applies wavelet-based feature extraction specifically for binary VPN detection.
+While both studies use wavelet decomposition, our work differs from VNAT in three key aspects: objectives (application classification vs.
+binary VPN detection), methodology (we systematically evaluate decomposition depths J=12 and J=5, finding J=12 superior despite diminishing returns), and architecture (demonstrating that RF achieves comparable performance to VNAT’s specialized Prototypical Network).
 In this work, we focus on binary VPN detection and re-evaluate the effectiveness of wavelet transform-based features for distinguishing VPN from non-VPN traffic.
@@
 II-B Flow Processing and Analysis
-We utilized NFStream [Aouini2022], a flexible network data analysis framework, to process the dataset’s PCAP files.
+We utilized NFStream [14], a flexible network data analysis framework, to process the dataset’s PCAP files.
 While VNAT provides both traffic traces in PCAP format and pre-extracted statistical summaries in HDF (h5) format, we chose NFStream for its ability to seamlessly integrate plugins, particularly for calculating the Wavelet Features discussed in Section II-C.
@@
 Wavelet transform is a time-frequency analysis method that decomposes signals into components at different scales, making it suitable for analyzing non-stationary traffic patterns.
-Unlike Fourier transforms, wavelets preserve both time and frequency information, capturing traffic bursts and periodic behaviors effectively [Ahad2016].
+Unlike Fourier transforms, wavelets preserve both time and frequency information, capturing traffic bursts and periodic behaviors effectively [15].
 The Discrete Wavelet Transform (DWT) decomposes a signal into approximation coefficients (low-frequency components) and detail coefficients (high-frequency components) at multiple levels.
@@
   4. 4.
-    Shannon Entropy ($H_{j}$) quantifies the information content [Shannon1948]:
+    Shannon Entropy ($H_{j}$) quantifies the information content [19]:
     $H_{j}=-\sum_{k}P_{j}[k]\log_{2}P_{j}[k]$ (6)
@@
 Our choice of decomposition level J=5 is supported by the packet count distribution in our filtered dataset, as shown in Table II.
-With the 25th percentile at 56 packets, J=5 (scale of $322^{5}=$ 32 according to the optimal decomposition level formula $J=\lfloor\log_{2}(n)\rfloor$) ensures effective capture of packet variations across the majority of flows.
-We also evaluate J=12 to enable direct comparison with [10044382], though at this level (scale of $40962^{12}=$ 4096), the decomposition scale significantly exceeds typical flow lengths in their dataset.
-The highly skewed distribution (mean $=2,599$, median $=63$) indicates that while some flows contain up to 857,900 packets, such extensive flows are rare outliers.
+With the 25th percentile at 56 packets, J=5 (scale of $2^{5}=32$ according to the optimal decomposition level formula $J=\lfloor\log_{2}(n)\rfloor$) ensures effective capture of packet variations across the majority of flows.
+We also evaluate J=12 to enable direct comparison with [11], though at this level (scale of $2^{12}=4096$), the decomposition scale significantly exceeds typical flow lengths in their dataset.
+The highly skewed distribution (mean =2,599=2,599, median =63=63) indicates that while some flows contain up to 857,900 packets, such extensive flows are rare outliers.
 This distribution suggests that while J=12 might capture finer-grained variations in exceptionally long flows, it might provide diminishing returns for typical traffic patterns.
@@
 This work utilizes the VNAT dataset [11] as the primary data source.
-All processed datasets derived from these traces—including extracted network flows, filtered datasets, and wavelet-based feature representations—along with analysis scripts, trained models, and performance metrics necessary to reproduce our results are available at [github-repo], ensuring full reproducibility of the machine learning experiments.
+All processed datasets derived from these traces—including extracted network flows, filtered datasets, and wavelet-based feature representations—along with analysis scripts, trained models, and performance metrics necessary to reproduce our results are available at [20], ensuring full reproducibility of the machine learning experiments.
 Acknowledgement
@@
 This work was also supported by project TKP2021-NVA-02 (financed under the TKP2021-NVA scheme), implemented with support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund.
-[…]